OpenStudy (anonymous):

GR.12 MATH: please help // If y-3, y, and 3y+4 are consecutive terms in a geometric sequence, determine the value(s) of y.

OpenStudy (anonymous):

\[\frac{y-3}{y}=\frac{y}{3y+4}\] because it's geometric sequence

OpenStudy (anonymous):

(y-3)(3y+4)=y^2 then solve this quadratic equation and both solutions will work for the problem.

OpenStudy (anonymous):

Oh my god! I got it! Thanks :) :)

OpenStudy (anonymous):

you are welcome. glad I was able to help.